# Vecchia Gaussian Processes: Probabilistic Properties, Minimax Rates and Methodological Developments

### v3

@@ Abstract
Gaussian Processes (GPs) are widely used to model dependencies in spatial statistics and machine learning; however, exact inference is computationally intractable, with a time complexity of $O(n^{3})$. Vecchia approximation allows scalable Bayesian inference of GPs by introducing sparsity in the spatial dependency structure characterized by a directed acyclic graph (DAG). Despite its practical popularity, the approach lacks rigorous theoretical foundations, and the choice of DAG structure remains an open question. In this paper, we systematically study Vecchia GPs as standalone stochastic processes, uncovering key probabilistic and statistical properties. For probabilistic properties, we prove that the conditional distributions of the Matérn GPs, as well as their Vecchia approximations, can be characterized by polynomial interpolations. This allows us to prove a series of results regarding small ball probabilities and Reproducing Kernel Hilbert Spaces (RKHSs) of Vecchia GPs. For the statistical methodology, we provide a principled guideline for selecting parent sets as norming sets with fixed cardinality, and we develop algorithms along these guidelines. In terms of theoretical guarantees, we establish posterior contraction rates for Vecchia GPs in the nonparametric regression model and show that minimax-optimal rates are attained under oracle rescaling or hierarchical Bayesian tuning. We demonstrate our theoretical results and methodology through numerical studies and provide an efficient implementation of our methods in C++, with interfaces in R.
62G08,
60G15, 62H11,
Gaussian Process,
Minimax Rate,
Norming Set,
@@ 1.1 Background and related work
Gaussian processes have seen a wide range of applications amongst others in spatial statistics (e.g., [21, 11, 20]), machine learning (e.g., [55, 34, 38]) and epidemiology [6] due to their flexibility in characterizing dependency structures and convenience for uncertainty quantification. However, exact statistical inference for Gaussian Processes (GPs), including maximum likelihood estimation of GP parameters and posterior inference based on GP priors, suffers from a computational complexity of $O(n^{3})$, due to the need to compute the determinant and inverse of the $n\times n$ covariance matrix. A significant body of work has focused on scalable inference methods for Gaussian Processes, including but not limited to inducing variable or reduced rank approximations [43, 10, 4, 50], covariance tapering [17, 28, 47], compositional likelihoods [2, 5], distributed methods [14, 13, 49] and Vecchia approximations. Vecchia approximations of Gaussian processes (Vecchia GPs), named after Aldo V. Vecchia [53], are defined as follows. First, let us decompose the density of a random vector $Z_{\mathcal{X}_{n}}=(Z_{X_{1}},...,Z_{X_{n}})\in\mathbb{R}^{n}$ into a product of univariate conditional densities,
$p(Z_{\mathcal{X}_{n}})=p(Z_{X_{1}})\prod_{i=2}^{n}p(Z_{X_{i}}|Z_{X_{1}},Z_{X_{2}},\cdots,Z_{X_{i-1}}).$
Note, that computing these conditional densities can be expensive. Hence, in Vecchia approximations each conditional set is replaced with its subset, known as the parent set, and the joint density is approximated as
$p(Z_{\mathcal{X}_{n}})\approx p(Z_{X_{1}})\prod_{i=2}^{n}p(Z_{X_{i}}|Z_{\mathrm{pa}(X_{i})}),$
@@
where $Z_{\mathrm{pa}(X_{i})}\subset\{Z_{X_{1}},...,Z_{X_{i-1}}\}$. Applying the above approximation to the finite dimensional marginals of a GP results in, under weak assumptions, a new GP [^1] For a proof, we refer to [12] or [56]., called the Vecchia GP. We provide a more formal description of the process in Section 2.2. The approximated GP is referred to as the mother GP. The underlying conditional dependence structure in the approximation, i.e. the set of parents for each covariate, is summarized by a DAG.
@@
There is an extensive literature about Vecchia approximations of GPs (e.g., [12, 27, 26, 42, 9]). Their popularity can mostly be attributed to two practical advantages. First, under the condition that all parent sets have bounded cardinality, Vecchia approximations possess $O(n)$ computational complexity for evaluating the joint density. This renders Veccha GPs as one of the best scalable methods for GP approximations. Second, the Vecchia approximation of a GP is a valid GP itself. Hence, inference via Vecchia GPs is statistically meaningful regardless of the quality of approximation of the mother GP.
@@
Despite their success in applications and popularity amongst researchers, there are major methodological and theoretical challenges that hinder further developments. On the methodological side, the optimal choice of the DAG structure is not clear, despite its vital role in the performance of Vecchia GPs. In fact, several of the proposed concepts even contradict each other. For example, many papers propose to choose close neighbors for parent sets [12, 57], while other authors promote remote locations [3] or random ordering [23]. Moreover, it is intuitively clear that larger parent sets lead to better approximations, but slower computations. However, there is no universally accepted proposal for the size of the parent set. In fact, the only qualitative results in the literature are of poly-log order [44, 25, 57], and it is not clear if further increasing the cardinality would result in non-marginal statistical advantages. We summarize the above challenges in the following two research questions:
@@
  • Q1 What is the minimal cardinality of the parent sets in the Vecchia GP to achieve optimal statistical inference?
  • Q2 How shall the parent sets be chosen to reach optimal inference?
@@
We answer these questions from a frequentist statistics perspective. We show that appropriately tuned Vecchia GPs can achieve optimal statistical inference in a minimax sense. More concretely, we investigate the posterior contraction rates of Vecchia GPs around the true functional parameter of interest and show that the minimax rates can be achieved for suitably chosen DAG structures and optimally tuned mother GPs.
@@
Despite the great popularity of Vecchia GPs in practice, their theoretical understanding were so far rather limited. In fact, we are only aware of the following few, recent theoretical papers. In the univariate case under certain assumptions on the variance, [56] derives asymptotic normality of Vecchia GPs. Under suitable conditions on the DAG structure, covariance function, and dataset, [44] shows that the precision matrix of a Vecchia GP closely approximates that of the corresponding mother GP. Additionally, [57] demonstrates that Vecchia GPs provide accurate approximations to the full GPs in Wasserstein distance. Furthermore, [25] derives that the maximum likelihood estimators of the Vecchia GP parameters has the same convergence and asymptotic normality properties as of the mother GPs.
@@
Although all the above papers offer solid contributions to the literature of Vecchia GPs, none of them take a frequentist perspective and address the contraction rate of the posterior around the truth in nonparametric models. In fact, all previous works tackle the theory of Vecchia GPs from the “approximation of the mother GP” perspective. They either directly quantify the approximation error between Vecchia GPs and the corresponding mother GPs, or study the statistical properties of the Vecchia GPs based on the approximation error bounds between the Vecchia GPs and their corresponding mother GPs. However, such indirect strategy might not lead to optimal result. Vecchia GPs are well-defined, practically relevant Gaussian processes themselves. Therefore their properties, both empirical and theoretical, shall be studied as standalone problems that do not require good approximations to their mother GPs as prerequisites. After all, the main goal is to provide a scalable process with similarly good statistical properties as the original GP. Restricting ourselves to an overly accurate approximation of the mother GP might prevent us to fully explore the computational advantages offered by the Vecchia approach.
@@
To investigate the statistical aspects of Vecchia GPs, one has to first understand its probabilistic properties, which so far were barely studied. More concretely, GPs are typically defined through their marginal distributions and the standard probabilistic techniques are tailored to such perspective, see [29, 36, 1]. However, Vecchia GPs are defined through the conditional distributions encoded in a DAG structure. For such constructions, to the best of our knowledge, no general tools are derived for computing the small deviation bounds and other probabilistic aspects of the GP. Hence, one of the main challenge of our work was to derive new mathematical techniques for studying the probabilistic properties of Vecchia approximated GPs. To highlight that the mother and Vecchia GPs can have substantially different properties, we note that in case of the popular Matérn process the mother GP is stationary while its Vecchia approximation is not anymore. In fact, little is known about the covariance kernel of the Vecchia GP or the corresponding Reproducing Kernel Hilbert Space (RKHS). Therefore, the techniques used for isotropic covariance functions, like the Matérn kernel, relying on spectral densities do not apply in case of Vecchia approximations anymore.
@@ 1.2 Our contributions
In this section, we collect our contributions on the probabilistic and statistical properties, and methodological developments of Vecchia Gaussian Processes. These results highly build on each other and extend out understanding of this approximation technique in several directions.
@@ 1.2.1 Probabilistic properties
Our contributions to the probabilistic aspects are twofold. First, we systematically study the conditional distributions that define the Vecchia GPs. We show that the conditional expectations of the Vecchia GPs can be characterized asymptotically with local polynomial interpolations on the parent sets. Furthermore, we prove that the conditional variances coincide with the approximation errors of these local polynomial approximations. Moreover, under suitable DAG structures, the degree of these local polynomials is exactly $\underline{\alpha}$, the largest integer strictly smaller than the regularity of the Matérn process. While the limit of Gaussian interpolation has been studied in the literature of radial basis functions [15, 32, 46, 35], to the best of our knowledge, we are the first to prove the stronger version with uniform convergence rate and to make this connection to the Vecchia GPs. On the other hand, the characterization of the variance with the approximation error is an entirely novel contribution.
@@
Our second contribution is the derivation of various probabilistic properties of Vecchia GPs based on the conditional probabilities instead of the standard route through marginal probabilities. One important quantity is the small deviation bound (i.e. small ball probability) of the GP, which is typically derived by its relation to the entropy of the unit ball of the associated RKHS [29, 36, 1, 22]. However, for Vecchia GPs the properties of the corresponding RKHS are not well characterized. In our work we provide a new set of tools, tailored to Vecchia GPs, but are potentially useful for other types of processes as well.
1.2.2 Nonparametric theory
@@
Building upon the probabilistic properties discussed above, we prove in context of the nonparametric regression model that appropriately tuned Vecchia GPs attain minimax posterior contraction rates for Hölder smooth functions. This demonstrates that Vecchia GPs can reduce the computational costs of their mother GPs considerably without sacrificing statistical accuracy. The optimal choice of the tuning/scaling parameter depends on the smoothness of the underlying deterministic functional parameter of interest, which is typically unknown in practice. Therefore, we also consider a data driven tuning of the GP, by endowing the scaling parameter with another layer of prior, forming a two-level, so called hierarchical prior distribution. We show that the posteriors arising from such hierarchical Vecchia GPs, under mild assumptions, concentrate their masses around the true function with the minimax rate.
1.2.3 Methodological developments
@@
Our research was partially motivated by the methodological questions Q1 and Q2 that remained open for a long time. We answer these questions from a frequentist perspective, focusing on optimal recovery of the underlying true function of interest in context of the nonparametric regression model. We show, that well chosen parent sets with cardinality equal to $\underline{\alpha}+d\choose\underline{\alpha}$} are sufficient to achieve minimax contraction rate with the Vecchia approximation. More concretely, we propose to choose a norming set as parents with a fixed, universal norming constant, see Section 3 for definition and background. For (approximately) grid data, we provide an explicit formulation of such sets, while for more general datasets we propose algorithms that can potentially find such norming sets given they exist. Numerical studies in Section 7 demonstrate that our method achieves almost identical performance to state-of-the-art methods for nonparametric function estimation, while substantially reducing the computational time.
@@
Our perspective and hence our answers to questions Q1 and Q2 are different from the bulk of the literature on Vecchia GPs. We believe that these differences reveal hidden features and characteristics and therefore contributes to the understanding of this popular, but so far theoretically not well underpinned method. In Section 8 we provide a detailed discussion of our results and their implications. Here we focus on one of the most important methodological aspect, i.e. the cardinality of the parent sets. The results in [44], [57] and [25] all require the parent sets to grow at poly-log rate in $n$, while we only need fixed sized parent sets. There is, however, no contradiction here. Previous works typically consider Vecchia GP approximations to the mother GPs and the poly-log size ensures vanishing approximation error. In contrast, we directly investigate the properties of the posterior resulting from the Vecchia GP as standalone prior and derive contraction rates in a nonparametric framework. Bypassing the approximation step and following such more direct approach, one can typically end up with weaker conditions.
1.3 Notations
@@
Next we introduce some notations used throughout the paper. For easier readability, we organize them into subsections.
@@ 1.3.1 Gaussian Processes
Let $\mathcal{X}\subset\mathbb{R}^{d}$ be the domain of the centered mother GP $(Z_{x},x\in\mathcal{X})$ with covariance kernel $K(\cdot,\cdot):\mathcal{X}\times\mathcal{X}\to\mathbb{R}$. We focus on the Matérn covariance function (Section 4.2.1 of [45]) with regularity hyper-parameter $\alpha>0$:
@@
where $\mathcal{K}_{\alpha}$ denotes the modified Bessel function of the second kind with parameter $\alpha$. It has a simple characterization in terms of Fourier transform
@@
where the constant in $\propto$” only dependends on $\alpha$ and $d$, see equation (11.4) of [22]. For $s,\tau>0$, the rescaled GP $(Z_{x}^{\tau,s},x\in\mathcal{X})$ is defined as
@@
We will refer to $\tau$ as time, while to $s$ as space (re-)scaling hyper-parameters.
Let us denote by $(\hat{Z}_{x}^{\tau,s},x\in\mathcal{X})$ the Vecchia approximation of the mother GP $(Z^{\tau,s}_{x},x\in\mathcal{X})$, as defined later in Section 2.2. When $\tau=s=1$, we abbreviate the notation as $(\hat{Z}_{x},x\in\mathcal{X})$ and let $\hat{K}(\cdot,\cdot):\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ denote the corresponding covariance kernel which is typically different from the mother covariance kernel $K(\cdot,\cdot)$. For arbitrary scaling parameters $\tau,s>0$ let $K^{\tau,s}(\cdot,\cdot)$ and $\hat{K}^{\tau,s}(\cdot,\cdot)$ denote the covariance functions of the corresponding rescaled processes. For notational convenience, we abbreviate the process $(Z^{\tau,s}_{x},x\in\mathcal{X})$ as $Z^{\tau,s}$. Similarly, let us introduce the abbreviations $\hat{Z}^{\tau,s}$, $Z$ and $\hat{Z}$. Furthermore, let $\mathbb{H}$ and $\mathbb{H}^{\tau,s}$ denote the RKHSs associated with the mother GPs $Z$ and $Z^{\tau,s}$, respectively. Similarly, let $\mathbb{H}_{n}$ and $\mathbb{H}_{n}^{\tau,s}$ be the RKHSs for the $n$-dimensional marginals on $\mathcal{X}_{n}$ of the processes $\hat{Z}$ and $\hat{Z}^{\tau,s}$, respectively.
For an arbitrary covariance function $K(\cdot,\cdot):\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$, and multi indices $k_{1},k_{2}\in\mathbb{N}^{d}$, let us denote by $K^{(k_{1},k_{2})}(x_{1},x_{2})=\frac{\partial^{k_{1}}}{\partial x_{1}^{k_{1}}}\frac{\partial^{k_{2}}}{\partial x_{2}^{k_{2}}}K(x_{1},x_{2})$ the $k_{1}$th and $k_{2}$th order derivatives with respect to the first and second arguments, respectively. Furthermore, for all finite sets $A,B\subset\mathcal{X}$, we denote the covariance matrix between $A$ and $B$ under the covariance function $K(\cdot,\cdot)$ as $K_{A,B}$. If $B=\{x\}$ contains a single element, we slightly abuse the notation by writing $K_{A,x}=K_{A,\{x\}}$.

### v4

# Supplementary Material to “Vecchia Gaussian processes: on probabilistic and statistical properties”

@@ Abstract
Gaussian Processes (GPs) are widely used to model dependencies in spatial statistics and machine learning. However, exact inference is computationally intractable for GP regression, with a time complexity of $O(n^{3})$. The Vecchia approximation scales up computation by introducing sparsity into the spatial dependency structure, represented by a directed acyclic graph (DAG). Despite its practical popularity, this approach lacks rigorous theoretical foundations, and the choice of DAG structure remains an open problem.
In this paper, we systematically study the Vecchia approximation of the popular, isotropic Matérn GP as standalone stochastic process and uncover key probabilistic and statistical properties. We propose selecting parent sets as norming sets with fixed cardinality in the Vecchia approximation. On the probabilistic side, we show that the conditional distributions of Matérn GPs, as well as their Vecchia approximations, can be characterized by polynomial interpolations. This enables us to establish several results on small ball probabilities and the Reproducing Kernel Hilbert Spaces (RKHSs) of Vecchia GPs. Building on these probabilistic results, we prove that in the nonparametric regression model, the corresponding posterior contracts around the truth at the optimal minimax rate, both under oracle rescaling and hierarchical tuning of the prior.
We illustrate the theoretical findings through numerical experiments on synthetic datasets. Our core algorithms are implemented in C++++ with an R interface.
@@ 1.1 Background and related work
Gaussian processes have seen a wide range of applications among others in spatial statistics (e.g., [21, 11, 20]), machine learning (e.g., [60, 35, 39]) and epidemiology [6] due to their flexibility in characterizing dependency structures and their convenience for uncertainty quantification. However, exact statistical inference in the Gaussian Process (GP) regression model – including maximum likelihood estimation of GP parameters and posterior inference based on GP priors – suffers from a computational complexity of $O(n^{3})$, owing to the need to compute the determinant and inverse of the $n\times n$ covariance matrix.
A substantial body of work has focused on scalable inference methods for Gaussian Processes, including but not limited to inducing variable or reduced rank approximations [44, 10, 4, 55], covariance tapering [17, 29, 50], compositional likelihoods [2, 5], distributed methods [14, 13, 53] and Vecchia approximations. Vecchia approximations of Gaussian processes (Vecchia GPs), named after Aldo V. Vecchia [58], are constructed as follows. First, consider decomposing the density of a random vector $Z_{\mathcal{X}_{n}}=(Z_{X_{1}},...,Z_{X_{n}})\in\mathbb{R}^{n}$ into a product of univariate conditional densities,
$p(Z_{\mathcal{X}_{n}})=p(Z_{X_{1}})\prod_{i=2}^{n}p(Z_{X_{i}}|Z_{X_{<i}}),$
where $Z_{X_{<i}}=\{Z_{X_{1}},...,Z_{X_{i-1}}\}$. The key idea behind the Vecchia approximation is to replace each conditioning set $Z_{X_{<i}}$ with a subset $Z_{\mathrm{pa}(X_{i})}\subseteq Z_{X_{<i}}$, often referred to as the parent set, thereby introducing sparsity into the implied graphical model. More formally, the joint density is approximated as
$p(Z_{\mathcal{X}_{n}})\approx p(Z_{X_{1}})\prod_{i=2}^{n}p(Z_{X_{i}}|Z_{\mathrm{pa}(X_{i})}).$
@@
Applying the above approximation to the finite dimensional marginals of a GP results, under mild assumptions, in a new GP, called the Vecchia GP. [^1] For a proof, we refer to [12] or [61]. We provide a more formal description of the process in Section 2.2. The approximated GP is referred to as the mother GP. The underlying conditional dependence structure encoded in the approximation – namely the set of parents for each covariate – is summarized by the DAG.
@@
There is a substantial literature on Vecchia approximations of GPs (e.g., [12, 28, 27, 43, 9]). Their popularity can largely be attributed to two practical advantages. First, when all parent sets have bounded cardinality, Vecchia approximations allow evaluation of the joint density in $O(n)$ computational time. This makes Vecchia GPs one of the most scalable methods for GP approximation. Second, the Vecchia approximation of a GP is itself a valid GP. Consequently, inference based on Vecchia GPs is statistically meaningful regardless of how well the approximation matches the mother GP.
@@
Despite their success in applications and widespread use in the literature, major methodological and theoretical challenges remain. Methodologically, the optimal choice of the DAG structure is unclear, even though it plays a crucial role in the performance of Vecchia GPs. Indeed, several proposed strategies contradict each other. For example, many papers advocate selecting nearby neighbors as parent sets [12, 62], while others propose choosing more distant locations [3] or even random ordering [23]. Moreover, while it is intuitively clear that larger parent sets yield better approximations at the expense of slower computation, there is no consensus on an appropriate parent-set size. The qualitative results available in the literature are only of polylogarithmic order [47, 26, 62], and it remains unclear whether further increasing the cardinality provides substantial statistical benefits. We summarize the above challenges in the following two research questions:
@@
  • Q1 What is the minimal cardinality of the parent sets in the Vecchia GP needed to achieve optimal statistical inference?
  • Q2 How should the parent sets be chosen to attain optimal inference?
@@
We address these questions from a frequentist perspective. We show that suitably tuned Vecchia GPs can achieve optimal statistical inference in a minimax sense. In particular, we investigate the posterior contraction rates of Vecchia GPs around the true functional parameter of interest and demonstrate that minimax-optimal rates can be achieved for appropriately chosen DAG structures and optimally tuned mother GPs.
@@
Despite the great popularity of Vecchia GPs in practice, their theoretical understanding has so far been rather limited. We are aware of only a few recent theoretical contributions. In the univariate case and under certain assumptions on the variance, [61] derives asymptotic normality of Vecchia GPs. Under suitable conditions on the DAG structure, covariance function, and dataset, [47] shows that the precision matrix of a Vecchia GP closely approximates that of the corresponding mother GP. Additionally, [62] demonstrates that Vecchia GPs provide accurate approximations to the full GPs in Wasserstein distance. Furthermore, [26] shows that the maximum likelihood estimators of the Vecchia GP parameters have the same convergence properties and asymptotic normality properties as those of the mother GP.
@@
Although all the above papers offer solid contributions to the literature of Vecchia GPs, none of them take a frequentist perspective, addressing the contraction rate of the posterior around the truth in nonparametric models. In fact, all previous works tackle the theory of Vecchia GPs from the perspective of “approximating the mother GP”. They either directly quantify the approximation error between Vecchia GPs and the corresponding mother GPs or study the statistical properties of Vecchia GPs based on error bounds between the two. However, such an indirect strategy might not lead to optimal results. Vecchia GPs are well-defined, practically relevant Gaussian processes in their own right. Therefore, their properties – both empirical and theoretical – shall be studied as standalone problems that do not require good approximations to their mother GPs as prerequisites. After all, the main goal is to provide a scalable process with statistical properties comparable to those of the original GP. Restricting ourselves to an overly accurate approximation of the mother GP might prevent us from fully exploring the computational advantages offered by the Vecchia approach.
@@
To investigate the statistical aspects of Vecchia GPs, one has to first understand their probabilistic properties, which so far have been barely studied. More concretely, GPs are typically defined through their marginal distributions and the standard probabilistic techniques are tailored to this perspective, see, for example [30, 37, 1]. However, Vecchia GPs are defined through conditional distributions encoded in a DAG structure. For such constructions, to the best of our knowledge, no general tools have been developed for computing small deviation bounds or other probabilistic characteristics of the GP. Hence, one of the main challenges of our work was to derive new mathematical techniques for studying the probabilistic properties of Vecchia approximated GPs. To highlight that the mother and Vecchia GPs can have substantially different properties, we note that in the case of the popular Matérn process, the mother GP is stationary, whereas its Vecchia approximation is not. In fact, little is known about the covariance kernel of the Vecchia GP or the corresponding Reproducing Kernel Hilbert Space (RKHS). Therefore, techniques used for isotropic covariance functions –such as the Matérn kernel – that rely on spectral densities no longer apply in the case of Vecchia approximations.
@@ 1.2 Our contributions
In this section, we collect our contributions on the probabilistic and statistical properties of Vecchia Gaussian Processes. We focus on isotropic Matérn processes, while extensions beyond this class of GPs are briefly discussed in Section 6. These results build heavily on one another and extend our understanding of this approximation technique in several directions.
@@ 1.2.1 Probabilistic properties
Our contributions to the probabilistic aspects are twofold. First, we systematically study the conditional distributions that define Vecchia GPs. We show that the conditional expectations of Vecchia Matérn GPs can be characterized asymptotically via local polynomial interpolations on the parent sets. Furthermore, we prove that the conditional variances coincide with the approximation errors of these local polynomial approximations. Moreover, under suitable DAG structures, the degree of these local polynomials is exactly $\underline{\alpha}$, the largest integer strictly smaller than the regularity of the Matérn process. While the limits of Gaussian interpolation have been studied in the literature on radial basis functions [15, 33, 49, 36], to the best of our knowledge, we are the first to prove the stronger version with uniform convergence rate and to make this connection to Vecchia GPs. In contrast, the characterization of the variance through the approximation error is an entirely novel contribution.
@@
Our second contribution is the derivation of various probabilistic properties of Vecchia GPs based on their conditional probabilities, rather than following the standard route through marginal probabilities. One important quantity is the small deviation bound (i.e. small ball probability) of the GP, which is typically derived using its relation to the entropy of the unit ball of the associated RKHS [30, 37, 1, 22]. However, for Vecchia GPs, the properties of the corresponding RKHS are not well characterized. In our work, we provide a new set of tools tailored to Vecchia GPs, which are also potentially useful for other types of processes.
1.2.2 Statistical properties
@@
Building upon the probabilistic properties discussed above, we prove – within the context of the nonparametric regression model – that appropriately tuned Vecchia GPs attain minimax posterior contraction rates for Hölder smooth functions. This demonstrates that Vecchia GPs can substantially reduce the computational costs of their mother GPs without sacrificing statistical accuracy. The optimal choice of the tuning/scaling parameter depends on the smoothness of the underlying function of interest, which is typically not available. Therefore, we also consider a data-driven tuning strategy by endowing the scaling parameter with another layer of prior, forming a two-level, hierarchical prior distribution. We show that the posteriors arising from such hierarchical Vecchia GPs, under mild assumptions, concentrate their masses around the true function at the minimax rate. Alternatively, one may plug-in the marginal maximum likelihood estimators of the hyperparameters into the Vecchia GP. The two approaches possess similar theoretical properties; see, for instance, [54, 45]. In the present paper, we focus on the former, hierarchical approach and leave the analysis of the latter for future work.
@@
Our research was partially motivated by the methodological questions Q1 and Q2. We address these questions from a frequentist perspective, focusing on optimal recovery of the underlying true function in the nonparametric regression model. We show, that well-chosen parent sets with cardinality equal to $\underline{\alpha}+d\choose\underline{\alpha}$} are sufficient to achieve minimax contraction rate with the Vecchia approximation. More concretely, we propose selecting a norming set as parents with a fixed, universal norming constant, see Section B for the definition and background. For (approximately) grid-structured data, we provide an explicit construction of such sets, while for more general datasets we propose algorithms that can potentially identify such norming sets when they exist. Numerical studies in Section A demonstrate that our proposal achieves high estimation accuracy while maintaining fast computation.
@@
Our perspective and consequently our answers to questions Q1 and Q2 differ from much of the existing literature on Vecchia GPs. We believe that these differences reveal previously hidden features and characteristics, thereby contributing to a deeper understanding of this popular, yet so far theoretically underdeveloped method. In Section 6, we provide a detailed discussion of our results and their implications. Here we focus on one of the most important methodological aspect: the cardinality of the parent sets. The results in [47], [62] and [26] all require the parent sets to grow at a poly-logarithmic rate in $n$, whereas we require only fixed sized parent sets. There is, however, no contradiction here. Previous work typically considers Vecchia GP approximations to the mother GPs and the poly-logarithmic growth ensures vanishing approximation error. In contrast, we investigate directly the properties of the posterior induced by the Vecchia GP as standalone prior, and we derive contraction rates in a nonparametric framework. By bypassing the approximation step and taking this more direct approach, one can typically obtain weaker conditions.
1.3 Notation
@@
Next we introduce some notation used throughout the paper. For easier readability, we organize them into subsections.
@@ 1.3.1 Gaussian processes
Let $\mathcal{X}\subset\mathbb{R}^{d}$ be the domain of the centered mother GP $(Z_{x},x\in\mathcal{X})$ with covariance kernel $K(\cdot,\cdot):\mathcal{X}\times\mathcal{X}\to\mathbb{R}$. We focus on the Matérn covariance function (Section 4.2.1 of [48]) with regularity hyperparameter $\alpha>0$:
@@
where $\mathcal{K}_{\alpha}$ denotes the modified Bessel function of the second kind with parameter $\alpha$. The Matérn kernel also admits a simple Fourier characterization:
@@
where the proportionality constant depends only on $\alpha$ and $d$, see equation (11.4) of [22]. For $s,\tau>0$, the rescaled GP $(Z_{x}^{\tau,s},x\in\mathcal{X})$ is defined by
@@
We will refer to $\tau$ and $s$ as the time and space (re)scaling hyperparameters, respectively.
Let $(\hat{Z}_{x}^{\tau,s},x\in\mathcal{X})$ denote the Vecchia approximation of the mother GP $(Z^{\tau,s}_{x},x\in\mathcal{X})$, defined formally in Section 2.2. When $\tau=s=1$, we abbreviate this as $(\hat{Z}_{x},x\in\mathcal{X})$ and let $\hat{K}(\cdot,\cdot):\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ denote its covariance kernel, which is typically different from the mother covariance kernel $K(\cdot,\cdot)$. For arbitrary scaling hyperparameters $\tau,s>0$, let $K^{\tau,s}(\cdot,\cdot)$ and $\hat{K}^{\tau,s}(\cdot,\cdot)$ denote the covariance kernels of the corresponding rescaled mother and Vecchia processes, respectively. For notational convenience, we abbreviate $(Z^{\tau,s}_{x},x\in\mathcal{X})$ as $Z^{\tau,s}$, and similarly write $\hat{Z}^{\tau,s}$, $Z$ and $\hat{Z}$. Let $\mathbb{H}$ and $\mathbb{H}^{\tau,s}$ denote the RKHSs associated with the mother GPs $Z$ and $Z^{\tau,s}$, respectively. Likewise, let $\mathbb{H}_{n}$ and $\mathbb{H}_{n}^{\tau,s}$ denote the RKHSs for the $n$-dimensional marginals on $\mathcal{X}_{n}$ of the processes $\hat{Z}$ and $\hat{Z}^{\tau,s}$.
For an arbitrary covariance function $K(\cdot,\cdot):\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$, and multi-indices $k_{1},k_{2}\in\mathbb{N}^{d}$, we denote by $K^{(k_{1},k_{2})}(x_{1},x_{2})=\frac{\partial^{k_{1}}}{\partial x_{1}^{k_{1}}}\frac{\partial^{k_{2}}}{\partial x_{2}^{k_{2}}}K(x_{1},x_{2})$ the mixed partial derivatives of order $k_{1}$ and $k_{2}$ with respect to the first and second arguments, respectively. For any finite sets $A,B\subset\mathcal{X}$, we denote by $K_{A,B}$ the covariance matrix between $A$ and $B$ under the covariance kernel $K(\cdot,\cdot)$. If $B=\{x\}$ is a singleton, we slightly abuse our notation and write $K_{A,x}=K_{A,\{x\}}$.